Three

times a​ number, minus 2
​,
is equal to two

times the​ number, plus 5
.
Find the number.

Answer by ikleyn(52786)

.

    3x - 2 = 2x + 5


    3x - 2x = 5 + 2


     x      = 7.              ANSWER